A radiator contains 15 quarts of fluid, 40% of which is antifreeze. How much fluid should be drained and replaced with pure antifreeze so that the new mixture is 60% antifreeze?
Answer by mananth(16946)

percent ---------------- quantity
Antifreeze I 40 ---------------- 5 quarts
Antifreeze II 100 ---------------- x quarts
Mixture 60 ---------------- 15 + x quarts

40 * 5 + 100 x = 60 ( 15 + x )
200 + 100 x = 900 + 60 x
100 x + -60 x = 900 - -200
40 x = 700
/ 40
x = 17.5
17.5 quarts of 100 % Antifreeze II
So 17.5 quarts has to be drained
m.ananth@hotmail.ca